Reports 1-1 of 1 Clear search Modify search
AdV-DAQ (Calibration)
Paoletti - 16:45 Friday 12 April 2024 (63949) Print this report
Hrec_hoft_20000Hz_Gated_500Hz no more available?
Images attached to this report
Comments to this report:
mours - 19:43 Friday 12 April 2024 (63952) Print this report

Indeed, the filter producing this channel was in an abnormal state. This was due to a too large change of the dynamic of the h(t) channel, due to weird h(t) values produced on April 8.

The filtering process has been restarted and the Hrec_hoft_20000Hz_Gated_500Hz channel is back.

The library version will be change during the Tuesday maintenance to use a more recent Fd version not sensitive to this kind of glitches.

Search Help